x/6+5=2x/3: x/6+5=2x/3
Answer by checkley77(3639)

x/6+5=2x/3
x/6-2x/3=-5
(x-2*2x)/6=-5
(x-4x)/6=-5
-3x/6=-5
-3x=6*-5
-3x=-30
x=-30/-3
x=10 answer.
proof:
10/6+5=2*10/3
(10+5*6)/6=20/3
40/6=20/3
120=120